OpenStudy (anonymous):

I need to solve this equation: Tg(x+45)=1+sen2x

OpenStudy (anonymous):

what is mean of sen??

OpenStudy (helder_edwin):

it is two equations! you typed something wrong!

OpenStudy (anonymous):

sen is the trigonometry function, seno in spanish

OpenStudy (anonymous):

Did somebody solve it? I tried and the results are \[x=\pi+2k \pi x=2\pi+2k \pi ?\]

OpenStudy (anonymous):

For English speakers, sen is sin (aka sine).

OpenStudy (anonymous):

normally tg 45 and tg (pi)+45=1 so x can be 0 and (pi)

OpenStudy (anonymous):

thanks!!

OpenStudy (anonymous):

and I have another one: \[\sin (\pi/6+x)=\sin (3\pi/4)\]

OpenStudy (anonymous):

(\[\Pi/6\]+x) can be \[3\Pi/4\] and \[\Pi/4\] so x can be \[\Pi/12 and 7\Pi/12\]

OpenStudy (anonymous):

would you mind telling me why you say that \[\pi/6+x\] can be \[3\pi/4\]? I do not see it...

OpenStudy (anonymous):

I am sorry I type it wrong!!! Sorry!! It is = \[\sin (3\pi/4 +2x)\]

OpenStudy (anonymous):

sin((pi)/6+x)=sin(3(pi)/4) degrees must be equal so that ((pi)/6+x) must be equal to the (3(pi)/4)

OpenStudy (anonymous):

Did you see the correction, sorry!! I

OpenStudy (anonymous):

I think you can solve this problem, if you can't i'll help you :)

OpenStudy (anonymous):

:) I let you know!!

OpenStudy (anonymous):

It must be the same idea that \[x+\pi/6 = 3\pi/4 + 2x \] and then x = \[-7\pi/12\]

OpenStudy (anonymous):

Congratulations :) That's right answer ;)

OpenStudy (helder_edwin):

is your first equation \[ \Large \tan(x+45º)=1+\sin(2x) \] or \[ \Large \tan(x+45º)=1+\sin^2x \]

OpenStudy (anonymous):

1+ sin (2x)